OpenStudy (anonymous):

What fraction of a given sample of radioactive nuclide remains after four half lives??

OpenStudy (anonymous):

@kelly226

OpenStudy (anonymous):

1 half-life: 1/2 left 2: half of that, or 1/4 3: half of that, or 1/8 4: half of that, or 1/16 In general: (1/2)^x, where x is the number of half-lives.

jimthompson5910 (jim_thompson5910):

After one half life, half of the substance is gone. After another half life, half of the remaining amount is gone. So you have (1/2)*(1/2) = 1/4 of the amount left over After another half life, half of the remaining amount is gone. So you have (1/2)*(1/2)*(1/2) = 1/8 of the amount left over See where this is going?

OpenStudy (anonymous):

THANKS!